# Infrared finite scattering theory: Scattering states and representations of the BMS group

### v1

@@ Abstract
Any non-trivial scattering with any massless fields in four spacetime dimensions will generically produce an “out” state with memory which gives rise to infrared divergences in the standard $S$-matrix. To obtain an infrared-finite scattering theory, one must suitably include states with memory. However, except in the case of QED with massive charged particles, asymptotic states with memory that have finite energy and angular momentum have not been constructed for more general theories (e.g. massless QED, Yang-Mills and quantum gravity). To this end, we construct direct-integral representations over the “Lorentz orbit” of a given memory and classify all “orbit space representations” that have well-defined energy and angular momentum. We thereby provide an explicit construction of a large supply of physical states with memory as well as the explicit action of the BMS charges all states. The construction of such states is a key step toward the formulation of an infrared-finite scattering theory. While we primarily focus on the quantum gravitational case, we outline how the methods presented in this paper can be applied to obtain representations of the Poincaré group with memory for more general quantum field theories.
@@ 1 Introduction
In the context of collider physics — where the relevant scattering timescales and detector resolution are insensitive to low-frequency physics — one can impose an infrared cut-off and restrict attention to “inclusive cross-sections” for the scattering of any “hard” particles [11, 12, 13, 14]. Nevertheless, on longer time scales, the production of soft quanta cannot be ignored resulting in, for example, the complete decoherence [^1] In four-dimensional, flat spacetime, this decoherence rate due to soft quanta grows logarithmically in time [15]. However, in the presence of a black hole or cosmological horizon in any spacetime dimension, the analogous production of (massive or massless) soft quanta through the horizon results in a decoherence that grows linearly in time [16, 17, 18]. of any “hard” particles at late times [15, 19]. Furthermore, scattering theory as well as the asymptotically early and late time behavior of states play a central role in any holographic description of quantum gravity such as in celestial holography (see, e.g., [20, 21, 22, 23] and references therein). However, since scattering will generically produce memory, the standard $S$-matrix cannot be defined since there are no states with memory in the standard Fock space. Therefore, in order to keep track of all observables — including low-frequency observables such as the memory — it is essential to formulate a well-defined, infrared-finite scattering theory.
@@
There is, however, a major issue in carrying out this proposal — or any other proposed formulation of an infrared-finite scattering theory. As we have just explained, one does not require a “preferred” Hilbert space to formulate a well-defined, infrared-finite scattering theory. However, one does need to construct “sufficiently many” physical states with finite energy and angular momentum with non-vanishing memory. Given any in-state with finite energy and angular momentum, the out-state will generically be a state with (conserved) energy and angular momentum and non-zero memory. While in QED with massive charges, the Faddeev-Kulish construction provides us with a large class of physical states with memory, states in massless QED, Yang-Mills or quantum gravity with finite energy-momentum, angular momentum and non-vanishing memory have never been constructed.
@@
is finite. [^4] While it is tempting to restrict to smooth memories, instead of the bigger space of square-integrable ones, it will be argued in a forthcoming paper [42] that a generic scattering process necessarily produces non-smooth but square-integrable memory. A much more serious issue arises when one seeks a well-defined action the angular momentum. The memory is not invariant under general Lorentz transformations and, therefore, any eigenstates of memory in $\mathscr{F}_{\Delta}$ do not have well-defined Lorentz charges (see also [10]). Nevertheless, the space of memories obtained by acting on a given memory $\Delta_{AB}$ with the full Lorentz group is a finite-dimensional manifold — the orbit space $\mathscr{O}$ associated to the given memory $\Delta_{AB}$. If $y\in\mathscr{O}$ denotes a point in the orbit space, the action of the Lorentz group yields a collection Fock spaces $\mathscr{F}_{\Delta(y)}$ over each point in $\mathscr{O}$. Furthermore, each orbit space admits a unique, Lorentz invariant measure $\mu(y)$. Thus, associated to each memory $\Delta_{AB}$, we obtain a direct integral Hilbert space [^5] The direct integral given by eq. 1.3 is over a finite-dimensional subspace of memories. One could instead attempt to consider a direct integral over the full, infinite-dimensional space of memories. This case was considered in [43] and § 7 of [32] and requires a choice of infinite dimensional measure. While this choice is not unique, it was argued in [32] that there does not exist any Gaussian measure over the space of memories which yields a representation with well-defined energy and angular momentum.
@@
Finally, we compare and contrast the approach presented in this paper to other possible approaches in obtaining representations of the BMS group with memory. We first recall that there are at least two approaches to building the standard (zero memory) Fock space $\mathscr{F}_{0}$. One could either seek an irreducible representation of the field algebra which contains the invariant vacuum state or one could equivalently build $\mathscr{F}_{0}$ from the irreducible representations of the Poincaré group found by Wigner [45]. Thus, there is a clear compatibility between the irreducible representations of the quantum fields and the irreducible representations of the Poincaré group. This compatibility fails for massless fields. The irreducible representations of the quantum fields are the memory Fock representations $\mathscr{F}_{\Delta}$ which do not have a well-defined action of the BMS group (or any Poincaré subgroup). Furthermore, while irreducible representations of the BMS group were classified by McCarthy [46, 44, 47], it is not known to us whether any representation of the field algebra that admits a unitary representation of the (infinite dimensional) BMS group can be decomposed in terms of McCarthy’s irreducible representations. [^6] A theorem by Mautner [48] shows that any representation of a locally compact group can be written as a direct integral over irreducible representations. However, the BMS group is not locally compact, so one cannot argue that any representation is decomposable in terms of irreducible ones. A partial relationship between the representations of the asymptotic field algebra and McCarthy’s irreducible representations of the BMS group will be shown in [42]. Therefore, the approach taken in this paper is to consider representations of the full algebra of fields and BMS/Poincaré charges as opposed to seeking representations of any subalgebra.
@@ 1 Algebra of asymptotic quantum fields
We now describe the quantization of the radiative degrees of freedom at null infinity. The radiative degrees of freedom of the gravitational field are encoded in the shear tensor $\sigma_{AB}(x)$ on $\mathscr{I}$ where $x\in\mathscr{I}$. Here, and in the following, $u$ denotes the retarded/advanced time coordinate on $\mathscr{I}$ and $x^{A}$ are some choice of coordinates on $\mathbb{S}^{2}$ and so $x=(u,x^{A})$. The capital indices $A,B,\ldots$ are abtract indices of tensor fields on $\mathbb{S}^{2}$. The Bondi News tensor is defined as
@@ 2 Extension of the asymptotic algebra to include the BMS charges
$\delta_{(f,X)}s^{AB}$ $=\left(f+\tfrac{1}{2}\mathscr{D}_{C}X^{C}\right)\partial_{u}s^{AB}+\pounds_{X}s^{AB}+\tfrac{1}{2}(\mathscr{D}_{C}X^{C})s^{AB}$ (2.15)
@@
where $[X_{1},X_{2}]$ is the Lie bracket of $X_{1}^{A}$ and $X_{2}^{A}$. The memory observable has vanishing Poisson brackets the News, itself and the supertranslation charge
@@ 3 Memory Fock representations of the asymptotic field algebra
Then, extending this map so that it preserves the identity $\boldsymbol{1}$, products of operators and the ***-relation we get an automorphism of the subalgebra $\mathscr{A}_{\Delta}$ generated by $\boldsymbol{1}$, $\boldsymbol{N}(s)$ and $\boldsymbol{\Delta}(f)$. Using this automorphism, we can define a new state $\omega_{\nu}$ whose correlation functions are given by
@@
for any $\boldsymbol{O}\in\mathscr{A}_{\Delta}$. The state $\omega_{\nu}$ is a Gaussian state where it’s $n$-point functions are shifted by the classical solution $\nu_{AB}(x)$, in particular, it has a definite memory given by $\Delta(\nu,f)$. The GNS construction over this “shifted vacuum” i.e. acting with $\boldsymbol{N}(s)$ yields a Fock representation $\mathscr{F}_{\nu}$. If $\Delta(\nu,f)=0$ then the automorphism $\mathfrak{a}_{\nu}$ can be represented in the standard Fock space $\mathscr{F}_{0}$ by a unitary operator and the new state $\omega_{\nu}$ is a coherent state in $\mathscr{F}_{0}$. Other the other hand, if $\Delta(\nu,f)$ is chosen to be non-vanishing, then $\mathfrak{a}_{\nu}$ cannot be unitarily implemented in $\mathscr{F}_{0}$ and the new Fock space $\mathscr{F}_{\nu}$ is unitarily inequivalent to the standard Fock space $\mathscr{F}_{0}$. The choice of “vacuum” state $\omega_{\nu}$ of $\mathscr{F}_{\nu}$ is not unique and one could have chosen a different News tensor $\nu^{\prime}_{AB}$ with the same memory, i.e., $\Delta(\nu^{\prime},f)=\Delta(\nu,f)$ for all $f$. Nevertheless, as long as $\nu^{\prime}_{AB}$ is sufficiently regular [^11] Two Fock representations $\mathscr{F}_{\nu}$ and $\mathscr{F}_{\tilde{\nu}}$ with radiation fields $\nu_{AB}=2\partial_{u}\sigma_{AB}$ and $\tilde{\nu}_{AB}=2\partial_{u}\tilde{\sigma}_{AB}$ are unitarily equivalent if and only if the norm of $\sigma_{AB}-\tilde{\sigma}_{AB}$ (defined by eq. 2.20) is finite. Even if the radiation fields have the same memory, this norm may diverge if they are not sufficiently regular., then the two representations will be unitarily equivalent. Thus the unitary equivalence class of all $\mathscr{F}_{\nu}$ with sufficiently regular representative $\nu_{AB}$ with memory $\Delta_{AB}$ can be labeled by the memory of the representative — i.e. we may label the “memory representations” as $\mathscr{F}_{\Delta}$ rather than $\mathscr{F}_{\nu}$. Each memory Fock representation $\mathscr{F}_{\Delta}$ is unitarily inequivalent to $\mathscr{F}_{\Delta^{\prime}}$ for $\Delta\neq\Delta^{\prime}$.
@@ 1 General construction of representations of the BMS group with memory
To specify the action of the Lorentz group on $\mathscr{F}_{{\rm DI},\Delta}$ it will be useful to pick a family of News tensors $\nu_{AB}(x;y)$ and a corresponding choice of vacua |$\ket{\omega_{\nu(y)}}$ in every Fock space $\mathscr{F}_{\Delta(y)}$. While the family of News tensors is freely specifiable, it will be convenient to choose $\nu_{AB}(x;y)$ to depend smoothly on $y\in\mathscr{O}_{\Delta}$. This choice will allow us to straightforwardly define the unitary action of the Lorentz group. Any other choice of family of News tensors on $\mathscr{O}_{\Delta}$ will be related to this choice by a unitary map, and thus, we may consider such a family without loss of generality.
@@
Now, let $\Lambda\in{\tt SL}(2,\mathbb{C})$ be a given Lorentz transformation. We wish to define its action on states |$\ket{\Psi}$ specified by families of vectors of the form eq. 3.8. First we consider the action of $\Lambda$ on the chosen family of News tensors $\nu_{AB}(x;y)$; let $\nu^{\prime}_{AB}(x;y)$ be the family of News tensors obtained by acting on the original family $\nu_{AB}(x;y)$ by $\Lambda$. Then, the original News tensor at the point $\Lambda^{-1}y$ gets mapped to the transformed News tensor at the point $y$ for all $y\in\mathscr{O}_{\Delta}$. That is we have
@@
As the next step, we define how the test tensors in the expression eq. 3.8 transform under the action of $\Lambda$. Given the family of test tensors of the original state we define the transformation so that at the point $y$ the transformed test tensor is the Lorentz transformation by $\Lambda$ of the original test tensor at $\Lambda^{-1}y$, i.e.
@@
Putting together the transformations of both the families of test tensors and choice of vacua we have that under a Lorentz transformation by $\Lambda\in{\tt SL}(2,\mathbb{C})$, the state |$\ket{\Psi}$ specified by the family of vectors in eq. 3.8 transforms to |$\ket{\Psi^{\prime}}$ specified by the famliy of vectors
@@
$\qquad=\sum_{n=0}^{\infty}~{}\int\limits_{\mathscr{I}^{n}}d^{3}x_{1}\cdots d^{3}x_{n}~{}\bigg{[}-i\mathfrak{L}_{X}\psi^{A_{1}B_{1}\ldots A_{n}B_{n}}(x_{1},\ldots,x_{n};y)\boldsymbol{N}_{A_{1}B_{1}}(x_{1})\cdots\boldsymbol{N}_{A_{i}B_{i}}(x_{i})\ket{\omega_{\nu(y)}}$
@@
The first term arises from differentiating (with respect to $\epsilon$) the Lorentz transformation of the test tensors where $\mathfrak{L}$ is the Lie derivative acting on $\psi^{A_{1}B_{1}\ldots A_{n}B_{n}}(x_{1},\ldots,x_{n};y)$ viewed as a smooth function on $\mathscr{I}^{n}\times\mathscr{O}_{\Delta}$. Explicitly, the generator $X$ can be decomposed as a vector $Y$ on $\mathscr{O}_{\Delta}$ and a vector $X^{A}$ on $\mathscr{I}$ and so $\mathfrak{L}_{X}\psi$ can be expressed as
$\mathfrak{L}_{X}\psi^{A_{1}B_{1}\ldots A_{n}B_{n}}(x_{1},\ldots,x_{n};y)=\left[-\pounds_{Y}+\sum_{j=1}^{n}\left(\pounds_{X_{j}}+\tfrac{1}{2}(\mathscr{D}_{C}X_{j}^{C})\right)\right]\psi^{A_{1}B_{1}\ldots A_{n}B_{n}}(x_{1},\ldots,x_{n};y)$ (3.18)
where each $X_{j}^{A}$ is a copy of the vector field $X^{A}$ on each of the $\mathscr{I}$ on which the test tensor depends. The second term in eq. 3.17 arises from differentiating the Lorentz action on the vacua in eq. 3.15 with respect to $\epsilon$. Recall that $\Lambda^{\prime}(y)$ is an element of the little group ${\tt L}_{\Delta(y)}$ at every point $y\in\mathscr{O}_{\Delta}$ and therefore the action of the corresponding Lorentz charge is well-defined on each memory Fock space $\mathscr{F}_{\Delta(y)}$. The Lorentz charge obtained by differentiating the little group unitary $\boldsymbol{U}(\Lambda^{\prime}(y))$ is $\boldsymbol{\mathcal{Q}}_{i^{0}}(X^{\prime}(y))$ where the Lorentz generator $X^{\prime}(y)$ is given by (using eq. 3.12)

### v2

@@ Abstract
Any non-trivial scattering with massless fields in four spacetime dimensions will generically produce an “out” state with memory which gives rise to infrared divergences in the standard $S$-matrix. To obtain an infrared-finite scattering theory, one must suitably include states with memory. However, except in the case of QED with massive charged particles, asymptotic states with memory that have finite energy and angular momentum have not been constructed for more general theories (e.g. massless QED, Yang-Mills and quantum gravity). To this end, we construct direct-integral representations over the “Lorentz orbit” of a given memory and classify all “orbit space representations” that have well-defined energy and angular momentum. We thereby provide an explicit construction of a large supply of physical states with memory as well as the explicit action of the BMS charges all states. The construction of such states is a key step toward the formulation of an infrared-finite scattering theory. While we primarily focus on the quantum gravitational case, we outline how the methods presented in this paper can be applied to obtain representations of the Poincaré group with memory for more general quantum field theories.
@@ 1 Introduction
In the context of collider physics — where the relevant scattering timescales and detector resolution are insensitive to low-frequency physics — one can impose an infrared cut-off and restrict attention to “inclusive cross-sections” for the scattering of any “hard” particles [11, 12, 13, 14]. Nevertheless, on longer time scales, the production of soft quanta cannot be ignored resulting in, for example, the complete decoherence [^1] In four-dimensional, flat spacetime, this decoherence rate due to soft quanta grows logarithmically in time [15, 16]. However, in the presence of a black hole or cosmological horizon in any spacetime dimension, the analogous production of (massive or massless) soft quanta through the horizon results in a decoherence that grows linearly in time [17, 18, 19, 20]. of any “hard” particles at late times [15, 21]. Furthermore, scattering theory as well as the asymptotically early and late time behavior of states play a central role in any holographic description of quantum gravity such as in celestial holography (see, e.g., [22, 23, 24, 25] and references therein). However, since scattering will generically produce memory, the standard $S$-matrix cannot be defined since there are no states with memory in the standard Fock space. Therefore, in order to keep track of all observables — including low-frequency observables such as the memory — it is essential to formulate a well-defined, infrared-finite scattering theory.
@@
There is, however, a major issue in carrying out this proposal — or any other proposed formulation of an infrared-finite scattering theory. As we have just explained, one does not require a “preferred” Hilbert space to formulate a well-defined, infrared-finite scattering theory. However, one does need to construct “sufficiently many” physical states with finite energy and angular momentum with non-vanishing memory. Given any in-state with finite energy and angular momentum, the out-state will generically be a state with finite energy, angular momentum and non-zero memory. While in QED with massive charges, the Faddeev-Kulish construction provides us with a large class of physical states with memory, states in massless QED, Yang-Mills or quantum gravity with finite energy-momentum, angular momentum and non-vanishing memory have never been constructed.
This problem is a practical one and is well-illustrated by considering the example of the scattering of two gravitons. In this case, the “in” state is a state in the standard Fock space with finite energy and angular momentum. However, the “out” state will consist of outgoing “hard” gravitons together with soft radiation with memory. The precise distribution of memory depends strongly on the angular distribution of the incoming and outgoing gravitons. [^4] This point is more precisely explained in [44]. For different incoming scattering angles, the “out” state will generally have different distributions in memory and therefore will lie in inequivalent Hilbert space representations. The purpose of this paper is to construct a large supply of physical asymptotic states with finite energy, angular momentum and memory. We will do so by constructing an uncountably infinite number of unitarily inequivalent Hilbert spaces with memory that have a well-defined action of the BMS group. As we have already emphasized, the scattering theory cannot be formulated with a preferred (separable) Hilbert space and so we do not advocate attempting to construct an $S$-matrix using these representations. The scattering theory can be formulated in the absence of a preferred Hilbert space and the purpose is to characterize a large class of states which arise in scattering.
@@
is finite. [^5] While it is tempting to restrict to smooth memories, instead of the bigger space of square-integrable ones, it is shown in [44] that a generic scattering process necessarily produces non-smooth but square-integrable memory. A much more serious issue arises when one seeks a well-defined action of the angular momentum. The memory is not invariant under general Lorentz transformations and, therefore, any eigenstates of memory in $\mathscr{F}_{\Delta}$ do not have well-defined Lorentz charges (see also [10]). Nevertheless, the space of memories obtained by acting on a given memory $\Delta_{AB}$ with the full Lorentz group is a finite-dimensional manifold — the orbit space $\mathscr{O}$ associated to the given memory $\Delta_{AB}$. If $y\in\mathscr{O}$ denotes a point in the orbit space, the action of the Lorentz group yields a collection Fock spaces $\mathscr{F}_{\Delta(y)}$ over each point in $\mathscr{O}$. Furthermore, each orbit space admits a unique, Lorentz invariant measure $\mu(y)$. Thus, associated to each memory $\Delta_{AB}$, we obtain a direct integral Hilbert space [^6] The direct integral given by eq. 1.3 is over a finite-dimensional subspace of memories. One could instead attempt to consider a direct integral over the full, infinite-dimensional space of memories. This case was considered in [45] and § 7 of [34] and requires a choice of infinite dimensional measure. While this choice is not unique, it was argued in [34] that there does not exist any Gaussian measure over the space of memories which yields a representation with well-defined energy and angular momentum.
@@
Finally, we compare and contrast the approach presented in this paper to other possible approaches in obtaining representations of the BMS group with memory. We first recall that there are at least two approaches to building the standard (zero memory) Fock space $\mathscr{F}_{0}$. One could either seek an irreducible representation of the field algebra which contains the invariant vacuum state or one could equivalently build $\mathscr{F}_{0}$ from the irreducible representations of the Poincaré group found by Wigner [47]. Thus, there is a clear compatibility between the irreducible representations of the quantum fields and the irreducible representations of the Poincaré group. This compatibility fails for massless fields. The irreducible representations of the quantum fields are the memory Fock representations $\mathscr{F}_{\Delta}$ which do not have a well-defined action of the BMS group (or any Poincaré subgroup). Furthermore, while irreducible representations of the BMS group were classified by McCarthy [48, 46, 49], it is not known to us whether any representation of the field algebra that admits a unitary representation of the (infinite dimensional) BMS group can be decomposed in terms of McCarthy’s irreducible representations. [^7] A theorem by Mautner [50] shows that any representation of a locally compact group can be written as a direct integral over irreducible representations. However, the BMS group is not locally compact, so one cannot argue that any representation is decomposable in terms of irreducible ones. A partial relationship between the representations of the asymptotic field algebra and McCarthy’s irreducible representations of the BMS group is shown in [44]. Therefore, the approach taken in this paper is to consider representations of the full algebra of fields and BMS/Poincaré charges as opposed to seeking representations of any subalgebra.
@@ 1 Algebra of asymptotic quantum fields
We now describe the quantization of the radiative degrees of freedom at null infinity. The radiative degrees of freedom of the gravitational field are encoded in the shear tensor $\sigma_{AB}(x)$ on $\mathscr{I}$ where $x\in\mathscr{I}$. Here, and in the following, $u$ denotes the retarded/advanced time coordinate on $\mathscr{I}$ and $x^{A}$ are some choice of coordinates on $\mathbb{S}^{2}$ and so $x=(u,x^{A})$. The capital indices $A,B,\ldots$ are abstract indices of tensor fields on $\mathbb{S}^{2}$. The Bondi News tensor is defined as
@@ 2 Extension of the asymptotic algebra to include the BMS charges
$\delta_{(f,X)}s^{AB}$ $=\left(f+\tfrac{1}{2}u\mathscr{D}_{C}X^{C}\right)\partial_{u}s^{AB}+\pounds_{X}s^{AB}+\tfrac{1}{2}(\mathscr{D}_{C}X^{C})s^{AB}$ (2.15)
@@
where $[X_{1},X_{2}]$ is the Lie bracket of $X_{1}^{A}$ and $X_{2}^{A}$. The memory observable has vanishing Poisson brackets with the News, itself and the supertranslation charge
@@ 3 Memory Fock representations of the asymptotic field algebra
Then, extending this map so that it preserves the identity $\boldsymbol{1}$, products of operators and the *-relation we get an automorphism of the subalgebra $\mathscr{A}_{\Delta}$ generated by $\boldsymbol{1}$, $\boldsymbol{N}(s)$ and $\boldsymbol{\Delta}(f)$. Using this automorphism, we can define a new state $\omega_{\nu}$ whose correlation functions are given by
@@
for any $\boldsymbol{O}\in\mathscr{A}_{\Delta}$. The state $\omega_{\nu}$ is a Gaussian state where its $n$-point functions are shifted by the classical solution $\nu_{AB}(x)$, in particular, it has a definite memory given by $\Delta(\nu,f)$. The GNS construction over this “shifted vacuum” i.e. acting with $\boldsymbol{N}(s)$ yields a Fock representation $\mathscr{F}_{\nu}$. If $\Delta(\nu,f)=0$ then the automorphism $\mathfrak{a}_{\nu}$ can be represented in the standard Fock space $\mathscr{F}_{0}$ by a unitary operator and the new state $\omega_{\nu}$ is a coherent state in $\mathscr{F}_{0}$. Other the other hand, if $\Delta(\nu,f)$ is chosen to be non-vanishing, then $\mathfrak{a}_{\nu}$ cannot be unitarily implemented in $\mathscr{F}_{0}$ and the new Fock space $\mathscr{F}_{\nu}$ is unitarily inequivalent to the standard Fock space $\mathscr{F}_{0}$. The choice of “vacuum” state $\omega_{\nu}$ of $\mathscr{F}_{\nu}$ is not unique and one could have chosen a different News tensor $\nu^{\prime}_{AB}$ with the same memory, i.e., $\Delta(\nu^{\prime},f)=\Delta(\nu,f)$ for all $f$. Nevertheless, as long as $\nu^{\prime}_{AB}$ is sufficiently regular [^12] Two Fock representations $\mathscr{F}_{\nu}$ and $\mathscr{F}_{\tilde{\nu}}$ with radiation fields $\nu_{AB}=2\partial_{u}\sigma_{AB}$ and $\tilde{\nu}_{AB}=2\partial_{u}\tilde{\sigma}_{AB}$ are unitarily equivalent if and only if the norm of $\sigma_{AB}-\tilde{\sigma}_{AB}$ (defined by eq. 2.20) is finite. Even if the radiation fields have the same memory, this norm may diverge if they are not sufficiently regular., then the two representations will be unitarily equivalent. Thus the unitary equivalence class of all $\mathscr{F}_{\nu}$ with sufficiently regular representative $\nu_{AB}$ with memory $\Delta_{AB}$ can be labeled by the memory of the representative — i.e. we may label the “memory representations” as $\mathscr{F}_{\Delta}$ rather than $\mathscr{F}_{\nu}$. Each memory Fock representation $\mathscr{F}_{\Delta}$ is unitarily inequivalent to $\mathscr{F}_{\Delta^{\prime}}$ for $\Delta\neq\Delta^{\prime}$.
@@ 1 General construction of representations of the BMS group with memory
To specify the action of the Lorentz group on $\mathscr{F}_{{\rm DI},\Delta}$ it will be useful to pick a family of News tensors $\nu_{AB}(x;y)$ and a corresponding choice of vacua |$\ket{\omega_{\nu(y)}}$ in every Fock space $\mathscr{F}_{\Delta(y)}$. While the family of News tensors is freely specifiable, it will be convenient to choose $\nu_{AB}(x;y)$ to depend smoothly on $y\in\mathscr{O}_{\Delta}$. This choice will allow us to straightforwardly define the unitary action of the Lorentz group. Any other choice of a family of News tensors on $\mathscr{O}_{\Delta}$ will be related to this choice by a unitary map, and thus, we may consider such a family without loss of generality.
@@
Now, let $\Lambda\in{\tt SL}(2,\mathbb{C})$ be a given Lorentz transformation. We wish to define its action on states |$\ket{\Psi}$ specified by families of vectors of the form eq. 3.8. First, we consider the action of $\Lambda$ on the chosen family of News tensors $\nu_{AB}(x;y)$; let $\nu^{\prime}_{AB}(x;y)$ be the family of News tensors obtained by acting on the original family $\nu_{AB}(x;y)$ by $\Lambda$. Then, the original News tensor at the point $\Lambda^{-1}y$ gets mapped to the transformed News tensor at the point $y$ for all $y\in\mathscr{O}_{\Delta}$. That is we have
@@
As the next step, we define how the test tensors in the expression eq. 3.8 transform under the action of $\Lambda$. Given the family of test tensors of the original state, we define the transformation so that at the point $y$ the transformed test tensor is the Lorentz transformation by $\Lambda$ of the original test tensor at $\Lambda^{-1}y$, i.e.
@@
Putting together the transformations of both the families of test tensors and choice of vacua we have that under a Lorentz transformation by $\Lambda\in{\tt SL}(2,\mathbb{C})$, the state |$\ket{\Psi}$ specified by the family of vectors in eq. 3.8 transforms to |$\ket{\Psi^{\prime}}$ specified by the family of vectors
@@
$\qquad=\sum_{n=0}^{\infty}~{}\int\limits_{\mathscr{I}^{n}}d^{3}x_{1}\cdots d^{3}x_{n}~{}\bigg{[}-i\mathfrak{L}_{X}\psi^{A_{1}B_{1}\ldots A_{n}B_{n}}(x_{1},\ldots,x_{n};y)\boldsymbol{N}_{A_{1}B_{1}}(x_{1})\cdots\boldsymbol{N}_{A_{n}B_{n}}(x_{i})\ket{\omega_{\nu(y)}}$
@@
The first term arises from differentiating (with respect to $\epsilon$) the Lorentz transformation of the test tensors where $\mathfrak{L}$ is the Lie derivative acting on $\psi^{A_{1}B_{1}\ldots A_{n}B_{n}}(x_{1},\ldots,x_{n};y)$ viewed as a smooth function on $\mathscr{I}^{n}\times\mathscr{O}_{\Delta}$. Explicitly, the generator $X$ can be decomposed as a vector field $Y$ on $\mathscr{O}_{\Delta}$ and a vector field $X^{A}$ on $\mathscr{I}$ and so $\mathfrak{L}_{X}\psi$ can be expressed as
$\mathfrak{L}_{X}\psi^{A_{1}B_{1}\ldots A_{n}B_{n}}(x_{1},\ldots,x_{n};y)$ (3.18)
=[£$\qquad\qquad=\left[-\pounds_{Y}+\sum_{j=1}^{n}\left(\tfrac{1}{2}u_{j}\mathscr{D}_{C}X_{j}^{C}\frac{\partial}{\partial u_{j}}+\pounds_{X_{j}}+\tfrac{1}{2}\mathscr{D}_{C}X_{j}^{C}\right)\right]\psi^{A_{1}B_{1}\ldots A_{n}B_{n}}(x_{1},\ldots,x_{n};y)$
where each $X_{j}^{A}$ is a copy of the vector field $X^{A}$ on each of the $\mathscr{I}$ on which the test tensor depends. The second term in LABEL:eq:infLor arises from differentiating the Lorentz action on the vacua in eq. 3.15 with respect to $\epsilon$. Recall that $\Lambda^{\prime}(y)$ is an element of the little group ${\tt L}_{\Delta(y)}$ at every point $y\in\mathscr{O}_{\Delta}$ and therefore the action of the corresponding Lorentz charge is well-defined on each memory Fock space $\mathscr{F}_{\Delta(y)}$. The Lorentz charge obtained by differentiating the little group unitary $\boldsymbol{U}(\Lambda^{\prime}(y))$ is $\boldsymbol{\mathcal{Q}}_{i^{0}}(X^{\prime}(y))$ where the Lorentz generator $X^{\prime}(y)$ is given by (using eq. 3.11)